OpenStudy (anonymous):

Use the distributive property to simplify the expression below: 5t - 3(1 + t) I need help with thisssss I don't understand ittt

OpenStudy (turingtest):

- 3(1 + t)=-3-3t that was our distribution. Do you not see how that was done?

OpenStudy (anonymous):

The distributive property states that a(b+c)=ab+ac. Therefore, -3(1+t)=-3-3t

OpenStudy (anonymous):

Okay So you basically Multiply each by -3?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

Yuppers

OpenStudy (anonymous):

Okay thank youuu (:

OpenStudy (turingtest):

in general you multiply whats inside the parentheses by what's on the outside (the coefficient)

OpenStudy (precal):

5t-3(1+t) distribute -3 5t-3-3t combine like terms 5t-3t 2t-3